OpenStudy (anonymous):

4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

Solve for x right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Start by distributing\[4(x-1)=2(x+2)+6\]\[4x-4=2x+4+6\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now solve for x

OpenStudy (anonymous):

That means we want x to be by itself and the rest of the numbers to be on the other side. What should be the next step?

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[4x−4=2x+4+6 \]\[4x-2x−4=2x-2x+4+6 \]\[2x−4=4+6 \]

OpenStudy (anonymous):

Now that we have the x on the left hand side let's keep it there. Now we want to get rid of that -4 in the left hand side so we need to add 4 to both sides What do we get after that?

OpenStudy (anonymous):

divid 2 from x

OpenStudy (anonymous):

Yes what's the answer then?

OpenStudy (anonymous):

x = 7 ?

OpenStudy (anonymous):

Yes good job!!